OpenStudy (anonymous):

Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = (x - 6)2 Shift the graph of y = x2 down 6 units. Shift the graph of y = x2 right 6 units. Shift the graph of y = x2 up 6 units. Shift the graph of y = x2 left 6 units.

OpenStudy (anonymous):

shifting up or down is a y-shift shifting left or right is an x-shift for y-shifts, shifting up is subtracting the value of the shift amount from y shifting down is adding the value of the shift amount to y

OpenStudy (anonymous):

for x-shifts, shifting right is subtracting the value of the shift amount from x shifting left is adding the value of the shift amount to x

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

@Algebraic! so the answer is Shift the graph of y = x2 up 6 units. right?

OpenStudy (anonymous):

your expression has a value subtracted from the x value

OpenStudy (anonymous):

do you see that?

OpenStudy (anonymous):

x-6

OpenStudy (anonymous):

so it corresponds to what sort of shift? (based on what I said for all the cases)

OpenStudy (anonymous):

@Algebraic! im still not sure I think thats what it is

OpenStudy (anonymous):

y shift up (C units) looks like: y-C = x^3 y shift down (C units) looks like: y+C =x^3 x shift right (C units) looks like: y=(x-C)^3 x shift left (C units) looks like: y=(x+C)^3

OpenStudy (anonymous):

I see so it would be that it shifts 6 units to the right @Algebraic!

OpenStudy (anonymous):

yes